# Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers

### 12 Jan 2024 v1

@@
Abstract.
Large language models have catalyzed an unprecedented wave in code generation. While achieving significant advances, they blur the distinctions between machine- and human-authored source code, causing integrity and authenticity issues of software artifacts. Previous methods such as DetectGPT have proven effective in discerning machine-generated texts, but they do not identify and harness the unique patterns of machine-generated code. Thus, its applicability falters when applied to code. In this paper, we carefully study the specific patterns that characterize machine and human-authored code. Through a rigorous analysis of code attributes such as length, lexical diversity, and naturalness, we expose unique patterns inherent to each source. We particularly notice that the structural segmentation of code is a critical factor in identifying its provenance. Based on our findings, we propose a novel machine-generated code detection method called DetectCodeGPT, which improves DetectGPT by capturing the distinct structural patterns of code. Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency. Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.
copyright: none
1. Introduction
The advent of large language models (LLMs) such as Codex (Chen et al., 2021) and ChatGPT (OpenAI, 2022) has revolutionized software engineering by automating program generation, once the exclusive domain of human intellect. LLMs trained on code (Allal et al., 2023; Nijkamp et al., 2023; Zheng et al., 2023; Rozière et al., 2023), sophisticated yet accessible, have democratized the act of coding, allowing individuals without formal expertise to effectively engage with software creation (Yetiştiren et al., 2023; Xu et al., 2022). Their ability to generate syntactically correct and functionally robust code challenges traditional dynamics and introduces a new era of efficiency and innovation in software creation, maintenance, and evolution.
@@ 1. Introduction
Recently, there has been a growing research trend in detecting machine-generated texts (Yang et al., 2023; Zhao et al., 2023). Perturbation-based methods like DetectGPT (Mitchell et al., 2023), have achieved state-of-the-art results in identifying machine-generated text. DetectGPT employs likelihood score discrepancies between the original text and its perturbed variants to enhance detection accuracy. And the perturbation process is performed by a pretrained language model (e.g., T5 (Raffel et al., 2020)) to mask portions of the original text and then reconstruct it. While natural language text can accommodate a degree of variation without losing coherence, code must adhere to rigid structural constraints. Perturbation strategies designed for text often lead to code syntax errors, thereby diminishing the efficacy of these methods when applied to machine-generated code (Lee et al., 2023).
In this paper, we conduct a comparative analysis of the patterns between machine and human-written code from five aspects, including length, lexical diversity, token frequency, token category proportions, and naturalness. Through our exploration, we uncovered that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces.
Capitalizing on our findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code. Our DetectCodeGPT improves DetectGPT by capturing the distinct structural patterns of code. Diverging from conventional techniques that depend on external pretrained model (e.g. CodeT5 (Wang et al., 2021)) to perform Masked Language Modeling (MLM) for perturbations, our DetectCodeGPT ingeniously perturbs the code corpus by strategically inserting spaces and newlines. Eschewing the need for an LLM for perturbations, our method provides a much higher efficiency, Since the proposed perturbation strategy well captures the structural distinctions between machine and human code, our DetectCodeGPT also ensures higher efficacy.
To rigorously evaluate the effectiveness of DetectCodeGPT, we conducted extensive experiments using two varied datasets across six code language models (LLMs), ranging from 1.3 to 7 billion parameters, and operating at two different decoding temperatures. The results demonstrate that, unlike in the text domain where training a supervised model is relatively straightforward (Yang et al., 2023; Tian et al., 2023), detecting machine-authored code presents a unique challenge. Nevertheless, DetectCodeGPT markedly outperforms contemporary state-of-the-art methods in the machine-generated code detection task, delivering a stable and significant improvement with an average absolute increase of 6.46% in AUROC.
@@
  • We propose a novel machine-authored code detection method called DetectCodeGPT leveraging the distinct structural patterns of code.
@@
2. Background
2.1. Large Language Models for Code
GPT language models (Radford et al., 2018, 2019) based on Transformer (Vaswani et al., 2017) decoder has achieved remarkable success in natural language processing tasks (Ray, 2023). In the file of code generation, Codex (Chen et al., 2021) and AlphaCode (Li et al., 2022) are pioneering works to train large language models on code. These models are usually trained with the MLE objective to maximize the likelihood of the next token given the previous tokens (Radford et al., 2018). Mathematically, this is captured as:
(1) $\theta^{*}=\arg\max_{\theta}\sum_{i}\log p_{\theta}(x_{i})$
where $x_{i}$ denotes each instance in the training dataset and $p_{\theta}(x_{i})$ represents the model’s probability distribution parameterized by $\theta$. And the training data often contain millions of code in different programming languages collected from open source repositories like GitHub (Husain et al., 2020; Kocetkov et al., 2022; Gao et al., 2020).
Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle (Wang et al., 2021; Fried et al., 2022; Nijkamp et al., 2023; Zheng et al., 2023) and also instruction fine-tuning (Luo et al., 2023; Rozière et al., 2023). And recent large language models pretrained on a mixture of code and natural language like ChatGPT (Schulman et al., 2022), GPT4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) has also shown promising results on code generation tasks benefiting from the enormous amount of data and model parameters.
2.2. Perturbation-Based Detection of Machine-Generated Text
In the realm of machine-generated text detection, perturbation based method like DetectGPT (Mitchell et al., 2023) stands as the state-of-the-art technology (Yang et al., 2023). In this section, we take DetectGPT as an example to illustrate the idea of perturbation-based detection methods. DetectGPT distinguishes between machine and human-generated text by analyzing the patterns in their probability scores (Mitchell et al., 2023). The core idea is that when a text $x$ generated from a machine is subtly changed to $\tilde{x}$ through a perturbation process $q(\cdot|x)$ (e.g., MLM with T5 (Ahmad et al., 2021)), there is a sharper decline in its log probability scores $\log p_{\theta}(x)$ than that in human-generated text. This is because a machine-generated text is usually more predictable and tightly bound to the patterns it was trained on, leading to a distinct negative curvature in log probability when the text is perturbed. By contrast, human-written texts are characterized by a rich diversity that reflects a blend of experiences and cognitive processes. As a result, it doesn’t follow such predictable patterns, and its log probability scores $\log p_{\theta}(x)$ do not plummet as dramatically when similarly perturbed. Based on such discrepancy, we can define a likelihood discrepancy score for each input code to measure the drop of log probability after perturbation.
(2) $\mathbf{d}\left(x,p_{\theta},q\right)\triangleq\log p_{\theta}(x)-\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log p_{\theta}(\tilde{x})$
@@ 2.2. Perturbation-Based Detection of Machine-Generated Text
3. Empirical Analysis
In this section, we conduct a comparative analysis of the distinct features of machine-generated and human-written code. We choose to assess code conciseness through length variability, evaluate vocabulary breadth via lexical diversity, understand common coding patterns by analyzing token frequency, identify stylistic differences through token category proportions, and gauge code fluency and predictability by measuring naturalness. These choices are aimed at providing a comprehensive analysis of coding styles between human and machine programmers.
3.1. Study Design
3.1.1. Length Variability
Code length stands as a cornerstone attribute when characterizing code (Barron et al., 1998; Albrecht and Gaffney, 1983; Rosenberg, 1997). The intricate balance of code length directly influences readability, maintainability, and even computational efficiency. A deeper grasp on how machine-generated code varies in length versus human-crafted code could reveal nuances about the inherent logic and preferences of LLMs, as well as the conventional coding patterns adopted by human developers. We investigate two metrics that characterize code length, namely, the number of tokens and the number of lines.
3.1.2. Lexical Diversity
Lexical diversity indicates the richness and variety of vocabulary present in a corpus. In the context of programming, this refers to the diversity in variable names, functions, classes, and reserved words. Analyzing lexical diversity offers a deeper understanding of the creativity, expressiveness, and potential complexity of code segments. There are two important empirical laws in both natural and programming languages revealing the lexical diversity: Zipf’s law (Zipf, 2016) and Heaps’ law (Heaps, 1978).
Zipf’s Law: It is initially identified in natural languages (Zipf, 2016), and later verified in the scope of programming languages (Zhang, 2008). It states that the frequency value $f$ of a token is inversely proportional to its frequency rank $r$: $f\propto\frac{1}{r^{\alpha}}$, where $\alpha$ is close to $1$ (Zipf, 2016). In programming languages, it states that a few variable names or functions are very commonly used across different scripts, while many others are rarely employed.
Heaps’ Law: Also verified in both text and code domain(Heaps, 1978; Zhang, 2009), Heaps’ Law characterizes the expansion of a vocabulary $V$ as a corpus $D$ increases in size: $V$ $\propto$ $D^{\beta}$, where $\beta$in (0, 1) captures the rate of vocabulary growth relative to the size of the corpus.
@@ 3.1.2. Lexical Diversity
3.1.3. Token Frequency
3.1.4. Token Category Proportion
Token category proportion refers to the proportion of tokens belonging to different categories (e.g., keyword, identifiers) in the code corpus. Understanding the proportion of token categories in code is akin to dissecting the anatomy of a language. It gives us a lens to view the nuances of coding style, the emphasis on structure, and the intricacies that distinguish human and machine-authored code.
To delve into the token category proportion, we analyze code with tree-sitter [^1] https://github.com/tree-sitter/tree-sitter and classify tokens into distinct categories, as detailed in Table 1. We then compute the proportion of each category in the code corpus.
Table 1. Studied categories of Python code tokens
@@ 3.1.4. Token Category Proportion
3.1.5. Naturalness
@@
The concept of code naturalness suggests that programming languages share a similar degree of regularity and predictability with natural languages (Hindle et al., 2016). This idea has been operationalized by employing language models to assess the probability of a specific token’s occurrence within a given context. Under this framework, we inspect how “natural” machine-generated code is compared to human-written code. We employ two metrics, likelihood and rank to measure the naturalness of each token in the studied code corpus. The likelihood stands for the probability $p$ of a token $x$ under the model $p_{\theta}$, denoted as $p_{\theta}(x)$. The rank of a token $x$ is the position of $x$ in the sorted list of all tokens based on their likelihoods, denoted as $r_{\theta}(x)$. Both metrics evaluate how likely a token is preferred by the model (Ippolito et al., 2020; Gehrmann et al., 2019; Solaiman et al., 2019). We calculate log scores on each token and then take the average to represent the whole code snippet as advised in (Gehrmann et al., 2019). To pinpoint the code elements that most significantly affect the score discrepancies, we also present the mean scores on different token categories in Table 1 for comparison.
3.2. Experimental Setup
@@
We choose the state-of-the-art CodeLlama model (Rozière et al., 2023) to generate code. Limited by our computational resources, we use the version with 7B parameters. As for the decoding strategies, we adopt the top-$p$ sampling method (Holtzman et al., 2020) with $p$=$0.95$ following (Chen et al., 2021). The temperature $T$ is an important parameter controlling the diversity of the generated code (Holtzman et al., 2020). Since current LLMs on code are usually evaluated across different decoding temperatures (Zheng et al., 2023; Nijkamp et al., 2022; Chen et al., 2021; Rozière et al., 2023), we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings. The maximum length of the generated code is set to 512 tokens. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
3.3. Dataset Preparation
@@
To compare with human-authored code, we extracted 10,000 Python functions randomly from the CodeSearchNet corpus (Husain et al., 2020), a compilation of open-source GitHub projects that provides a wide-ranging and varied codebase. We used function signatures and their accompanying comments as prompts for the model, and we collected the corresponding bodies of these functions to represent human-written code.
@@
While acknowledging that current models, including CodeLlama and even ChatGPT (Liu et al., 2023), may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet (Yetiştiren et al., 2023; Liu et al., 2023), the choice of this dataset is deliberate and insightful. Pitting the model against various real-world project code rather than simple programming problems, akin to those in the HumanEval (Chen et al., 2021) or MBPP (Austin et al., 2021) dataset, offers a more representative assessment. It allows us to analyze the differences between human and machine’s code when faced with broader, practical applications.
3.4. Results and Analysis
@@
3.4.1. Length Variability
(a) Token length ($T=0.2$)
(b) Token length ($T=1.0$)
(c) Line numbers ($T=0.2$)
(d) Line numbers ($T=1.0$)
Figure 1. Distribution of length for generated code under two different temperatures
One could hypothesize several reasons for these observed patterns. The propensity for conciseness at lower temperatures may reflect LLM’s training data, where probably concise solutions were more prevalent or deemed more ‘correct’. On the flip side, human developers, often juggling multiple considerations like future code extensions, comments for peer developers, or even personal coding style, might craft lengthier solutions. Furthermore, the narrowing of disparities at higher temperatures can be attributed to the model’s increased willingness to explore varied coding styles. At higher temperatures, the LLM possibly mimics a broader spectrum of human coding patterns, capturing the essence of diverse coding habits and styles found in its training corpus.
Finding 1: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their personal styles.
(a) Zipf’s Law ($T=0.2$)
(b) Zipf’s Law ($T=1.0$)
(c) Heaps’ Law ($T=0.2$)
(d) Heaps’ Law ($T=1.0$)
Figure 2. Comparison of lexical diversity on machine- and human-authored code
3.4.2. Lexical Diversity
Finding 2: Machines demonstrate a preference for a limited spectrum of frequently-used tokens, whereas human code exhibits a richer diversity in token selection.
3.4.3. Token Frequency
Table 2. Top 50 tokens from human- and machine-authored code from CodeLlama on CodeSearchNet with $T=0.2$

### 24 Jan 2024 v2

# Index Terms: Machine-generated code detection, large language models, code generation, empirical analysis

@@
Abstract
Large language models have catalyzed an unprecedented wave in code generation. While achieving significant advances, they blur the distinctions between machine- and human-authored source code, causing integrity and authenticity issues of software artifacts. Previous methods such as DetectGPT have proven effective in discerning machine-generated texts, but they do not identify and harness the unique patterns of machine-generated code. Thus, its applicability falters when applied to code. In this paper, we carefully study the specific patterns that characterize machine and human-authored code. Through a rigorous analysis of code attributes such as lexical diversity, conciseness, and naturalness, we expose unique patterns inherent to each source. We particularly notice that the syntactic segmentation of code is a critical factor in identifying its provenance. Based on our findings, we propose a novel machine-generated code detection method called DetectCodeGPT, which improves DetectGPT by capturing the distinct stylized patterns of code. Diverging from conventional techniques that depend on external LLMs for perturbations, DetectCodeGPT perturbs the code corpus by strategically inserting spaces and newlines, ensuring both efficacy and efficiency. Experiment results show that our approach significantly outperforms state-of-the-art techniques in detecting machine-generated code.
Index Terms:
Machine-generated code detection, large language models, code generation, empirical analysis
1 Introduction
The advent of large language models (LLMs) such as Codex [1] and ChatGPT [2] has revolutionized software engineering by automating program generation, once the exclusive domain of human intellect. LLMs trained on code [3, 4, 5, 6], sophisticated yet accessible, have democratized the act of coding, allowing individuals without formal expertise to effectively engage with software creation [7, 8]. Their ability to generate syntactically correct and functionally robust code challenges traditional dynamics and introduces a new era of efficiency and innovation in software creation, maintenance, and evolution.
@@ 1 Introduction
Recently, there has been a growing research trend in detecting machine-generated texts [9, 10]. Perturbation-based methods like DetectGPT [11], have achieved state-of-the-art results in identifying machine-generated text. DetectGPT employs likelihood score discrepancies between the original text and its perturbed variants to enhance detection accuracy. And the perturbation process is performed by a pretrained language model (e.g., T5 [12]) to mask portions of the original text and then reconstruct it. While natural language text can accommodate a degree of variation without losing coherence, code must adhere to rigid syntactic constraints. Perturbation strategies designed for text often lead to code syntax errors, thereby diminishing the efficacy of these methods when applied to machine-generated code [13].
In this paper, we conduct a comparative analysis of the patterns between machine- and human-authored code from three aspects, including lexical diversity, conciseness, and naturalness. Through our exploration, we uncovered that compared to human, machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespace tokens.
Capitalizing on our findings, we propose a novel method called DetectCodeGPT for detecting machine-authored code. Our DetectCodeGPT improves DetectGPT by capturing the distinct styles of code. Diverging from conventional techniques that depend on external pretrained model (e.g. CodeT5 [14]) to perform Masked Language Modeling (MLM) for perturbations, our DetectCodeGPT ingeniously perturbs the code by strategically inserting spaces and newlines. Eschewing the need for an LLM for perturbations, our method provides much higher efficiency. Since the proposed perturbation strategy well captures the stylistic distinctions between machine- and human-authored code, our DetectCodeGPT also ensures higher efficacy.
To rigorously evaluate the effectiveness of DetectCodeGPT, we conducted extensive experiments using two varied datasets across six code language models, ranging from 1.3 to 7 billion parameters, and operating at two different decoding conditions. The results demonstrate that, unlike in the text domain where training a supervised model is relatively straightforward [9, 15], detecting machine-authored code presents a unique challenge. Nevertheless, DetectCodeGPT markedly outperforms contemporary state-of-the-art methods in the machine-generated code detection task, delivering a stable and significant improvement.
@@
  • We propose a novel machine-authored code detection method called DetectCodeGPT leveraging the distinct stylistic patterns of code.
@@
2 Background
2.1 Large Language Models for Code
GPT language models [16, 17] based on Transformer [18] decoder has achieved remarkable success in natural language processing tasks [19]. In the file of code generation, Codex [1] and AlphaCode [20] are pioneering works to train large language models on code. These models are usually trained with the MLE objective to maximize the likelihood of the next token given the previous tokens [16]. Mathematically, this is captured as:
$\theta^{*}=\arg\max_{\theta}\sum_{i}\log p_{\theta}(x_{i})$ (1)
where $x_{i}$ denotes each instance in the training dataset and $p_{\theta}(x_{i})$ represents the model’s probability distribution parameterized by $\theta$. And the training data often contain millions of code in different programming languages collected from open source repositories like GitHub [21, 22, 23].
Later advances to improve LLMs on code include designing new pretraining tasks like fill-in-the-middle [14, 24, 4, 5] and also instruction fine-tuning [25, 6]. And recent large language models pretrained on a mixture of code and natural language like ChatGPT [26], GPT4 [27] and LLaMA [28] has also shown promising results on code generation tasks benefiting from the enormous amount of data and model parameters.
2.2 Perturbation-Based Detection of Machine-Generated Text
In the realm of machine-generated text detection, perturbation based method like DetectGPT [11] stands as the state-of-the-art technology [9]. In this section, we take DetectGPT as an example to illustrate the idea of perturbation-based detection methods. DetectGPT distinguishes between machine and human-generated text by analyzing the patterns in their probability scores [11]. The core idea is that when a text $x$ generated from a machine is subtly changed to $\tilde{x}$ through a perturbation process $q(\cdot|x)$ (e.g., MLM with T5 [29]), there is a sharper decline in its log probability scores $\log p_{\theta}(x)$ than that in human-generated text. This is because a machine-generated text is usually more predictable and tightly bound to the patterns it was trained on, leading to a distinct negative curvature in log probability when the text is perturbed. By contrast, human-written texts are characterized by a rich diversity that reflects a blend of experiences and cognitive processes. As a result, it doesn’t follow such predictable patterns, and its log probability scores $\log p_{\theta}(x)$ do not plummet as dramatically when similarly perturbed. Based on such discrepancy, we can define a likelihood discrepancy score for each input code to measure the drop of log probability after perturbation.
$\mathbf{d}\left(x,p_{\theta},q\right)\triangleq\log p_{\theta}(x)-\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log p_{\theta}(\tilde{x})$ (2)
@@ 2.2 Perturbation-Based Detection of Machine-Generated Text
3 Empirical Analysis
In this section, we conduct a comparative analysis of the distinct features of machine- and human-authored code.
3.1 Study Design
To gain insights into distinctions between human and machine programmers, we consider three primary aspects that are relevant to coding styles, namely diversity, conciseness, and naturalness [30, 31, 32, 33, 34], which can be measured by specific metrics.
3.1.1 Lexical Diversity
Lexical diversity indicates the richness and variety of vocabulary present in a corpus. In the context of programming, this refers to the diversity in variable names, functions, classes, and reserved words. Analyzing lexical diversity offers a deeper understanding of the creativity, expressiveness, and potential complexity of code segments. There are four important empirical metrics in both natural and programming languages revealing the lexical diversity: token frequency, syntax element distribution, Zipf’s law [35] and Heaps’ law [36].
Syntax Element Distribution refers to the proportion of syntax elements (e.g., keywords, identifiers) in the code corpus. Understanding the distribution of syntax elements in code is akin to dissecting the anatomy of a language. It gives us a lens to view the nuances of coding style, the emphasis on structure, and the intricacies that distinguish human- and machine-authored code.
To delve into the syntax element distribution, we analyze code with tree-sitter [^1] https://github.com/tree-sitter/tree-sitter and classify tokens into distinct categories, as detailed in Table I. We then compute the proportion of each category in the code corpus.
Zipf’s and Heaps’ Laws were initially identified in natural languages [35, 36], and later verified in the scope of programming languages [30, 31]. Zipf’s law states that the frequency value $f$ of a token is inversely proportional to its frequency rank $r$: $f\propto\frac{1}{r^{\alpha}}$, where $\alpha$ is close to $1$ [35]. In programming languages, it states that a few variable names or functions are very commonly used across different scripts, while many others are rarely employed. Heaps’ Law characterizes the expansion of a vocabulary $V$ as a corpus $D$ increases in size: $V$ $\propto$ $D^{\beta}$, where $\beta$in (0, 1) captures the rate of vocabulary growth relative to the size of the corpus.
@@ 3.1.1 Lexical Diversity
Table I: Studied categories of Python code tokens
@@
3.1.2 Conciseness
Conciseness stands as a cornerstone attribute when characterizing code [37, 33, 34]. The intricate balance of code conciseness directly influences readability, maintainability, and even computational efficiency. We investigate two metrics that characterize code conciseness, namely, the number of tokens and the number of lines.
Number of tokens gives us an indication of verbosity and complexity, showing the detailed composition of the code [37].
Number of lines helps us understand organizational choices, as spreading code across more lines can reflect a focus on readability and structure [34].
3.1.3 Naturalness
@@
The concept of code naturalness suggests that programming languages share a similar degree of regularity and predictability with natural languages [32]. This idea has been operationalized by employing language models to assess the probability of a specific token’s occurrence within a given context. Under this framework, we inspect how “natural” machine-generated code is compared to human-written code.
Token Likelihood and Rank are two metrics that measure the naturalness of each token in the studied code corpus. The token likelihood stands for the probability $p$ of a token $x$ under the model $p_{\theta}$, denoted as $p_{\theta}(x)$. The rank of a token $x$ is the position of $x$ in the sorted list of all tokens based on their likelihoods, denoted as $r_{\theta}(x)$. Both metrics evaluate how likely a token is preferred by the model [38, 39, 40]. We calculate log scores on each token and then take the average to represent the whole code snippet as advised in [39]. To pinpoint the code elements that most significantly affect the score discrepancies, we also present the mean scores on different syntax element categories in Table I for comparison.
3.2 Experimental Setup
@@
We choose the state-of-the-art CodeLlama model [6] to generate code. Limited by our computational resources, we use the version with 7B parameters. As for the decoding strategies, we adopt the top-$p$ sampling method [41] with $p$=$0.95$ following [1]. The temperature $T$ is an important parameter controlling the diversity of the generated code [41]. Since current LLMs on code are usually evaluated across different decoding temperatures [5, 42, 1, 6], we generate code with $T=0.2$ and $T=1.0$ to capture the model’s behavior under different settings. The maximum length of the generated code is set to 512 tokens. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
3.3 Dataset Preparation
@@
To compare with human-authored code, we extracted 10,000 Python functions randomly from the CodeSearchNet corpus [21], a compilation of open-source GitHub projects that provides a wide-ranging and varied codebase. We used function signatures and their accompanying comments as prompts for the model as in [1], and we collected the corresponding bodies of these functions to represent human-written code.
@@
While acknowledging that current models, including CodeLlama and even ChatGPT [43], may not yet craft code of unparalleled quality for intricate tasks such as those in CodeSearchNet [7, 43], the choice of this dataset is deliberate and insightful. Pitting the model against various real-world project code rather than simple programming problems, akin to those in the HumanEval [1] or MBPP [44] dataset, offers a more representative assessment. It allows us to analyze the differences between human and machine’s code when faced with broader, practical applications.
3.4 Results and Analysis
@@
3.4.1 Token Frequency
Table II: Top 50 tokens from human- and machine-authored code from CodeLlama